A 10.0 N melon is accelerated uniformly from rest at a rate of 2.50 m/s2. What is the kinetic energy of this object after it has accelerated at a distance of 15.0 m?
Answer #1

As you know weight of the melon

W=mg =10

m=10/g =10/9.8 =1.02 kg

Using equation of the motion

V^2=u^2+2aS

Here u=0, a=2.5m/s^2 , S=15 m

V^2=0+2×2.5×15

V=\sqrt{75} m/s

Kinetic energy K=mV^2/2

K=1.02×75/2

K=38.25 J. Answer